Three consecutive integers have a sum of 240. Find the integers.
Found 2 solutions by whty89, Paul:
Answer by whty89(20)

you know they are consecutive so divide 240 by 3
240/3=80 That is your middle number
to find other numbers just add 1 to 80, then subtract one from 80
The integers are 79,80,81

Answer by Paul(988)

That answer that is posted in wrong
3 consective integers are x, x+1, and x+2.
x+x+1+x+2=240
3x+3=240
3x=237
x=79
79+1=80
79+2=81

hence the 3 integers are 79, 80, and 81.
Paul.